# Bayesian Inference for Relational Graph in a Causal Vector Autoregressive Time Series

### v1

@@ A.2 Test construction
Let $\rho_{1},\ldots,\rho_{p}$ stand for the eigenvalues of $\bm{\Upsilon}_{T,0}^{-1/2}(\bm{\Upsilon}_{T,1}-\bm{\Upsilon}_{T,0})\bm{\Upsilon}_{T,0}^{-1/2}$. By Lemma 4, the Reyni divergence is given by $\frac{1}{4}\sum_{j=1}^{T}[2\log(1+\rho_{j})-\log(1+\rho_{j})]$. By arguing as in the proof of Lemma 1 of Roy et al. [2024], each term inside the sum can be bounded below by $c_{2}\min(\rho_{j}^{2},c_{1})$ for some constants $c_{1},c_{2}>0$, where $c_{1}$ can be chosen as large as we like at the expense of making $c_{2}$ appropriately smaller. Under Condition (A4), using Relation (iii) of Lemma 5, we obtain that the $\rho_{j}$ are bounded by some constant not changing with $T$. Thus, with a sufficiently large $c_{1}$, the minimum operation in the estimate is redundant; that is, the Reyni divergence is bounded below by $\frac{1}{4}c_{2}\|\bm{\Upsilon}_{T,0}^{-1/2}(\bm{\Upsilon}_{T,1}-\bm{\Upsilon}_{T,0})\bm{\Upsilon}_{T,0}^{-1/2}\|_{\mathrm{F}}^{2}\geq c_{2}^{\prime}T\epsilon_{T}^{2}$ for some constant $c_{2}^{\prime}>0$. This follows since $\|\bm{\Upsilon}_{T,0}^{-1/2}(\bm{\Upsilon}_{T,1}-\bm{\Upsilon}_{T,0})\bm{\Upsilon}_{T,0}^{-1/2}\|_{\mathrm{F}}\geq\|\bm{\Upsilon}_{T,1}-\bm{\Upsilon}_{T,0}\|_{\mathrm{F}}/\|\bm{\Upsilon}_{T,0}\|^{2}_{\mathrm{op}}$ and by Relation (iii) of Lemma 5, $\|\bm{\Upsilon}_{T,0}\|^{2}_{\mathrm{op}}\leq\|\bm{\Omega}^{-1}_{0}\|^{2}_{\mathrm{op}}$, which is bounded by a constant by Condition (A5). Thus both error probabilities are bounded by $e^{-c_{2}^{\prime}T\epsilon_{T}^{2}}$ for some $c_{2}^{\prime}>0$.
@@
${\mathbb{E}}_{\bm{\zeta}_{2}}(1-\phi_{T})$ $\leq[{\mathbb{E}}_{\bm{\zeta}_{1}}(1-\phi_{T})]^{1/2}[{\mathbb{E}}_{\bm{\zeta}_{1}}(Q_{\bm{\zeta}_{2}}/Q_{\bm{\zeta}_{2}})^{2}]^{1/2}$
@@
which is bounded by $\|\bm{f}_{1}\|_{\infty}+pdr\max_{j}(\|\bm{L}_{j,1}\|_{\infty}^{2})\leq C({T}\bar{\epsilon}_{T}^{2})^{1/\gamma}+pdr(T\bar{\epsilon}_{T}^{2})^{2/\gamma}\leq C^{\prime}dr(T\bar{\epsilon}_{T}^{2})^{2/\gamma}$ for some constant $C^{\prime}>0$ as $p$ is fixed and $T\bar{\epsilon}_{T}^{2}\to\infty$. Thus the expression in (A.2) is bounded by $\exp[-c_{2}T\epsilon_{T}^{2}+(C^{\prime}dr)^{2}(T\bar{\epsilon}_{T}^{2})^{4/\gamma}\delta_{T}^{2}]\leq\exp[-c_{3}T\epsilon_{T}^{2}]$ for some constant $c_{3}>0$ by choosing $\delta_{T}$ a sufficiently small constant multiple of $(C^{\prime}dr)^{-1}(T\bar{\epsilon}_{T}^{2})^{1/2-2/\gamma}$ and $M>0$ large enough.
@@
when $\{\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{F}},\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{F}},\|\bm{L}_{1,k}-\bm{L}_{2,k}\|^{2}_{\mathrm{F}}\}\lesssim\zeta_{T}^{2}$
We have, 1) $\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{F}}\leq dr\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\infty}$, 2) $\|\bm{L}_{1,j}-\bm{L}_{2,j}\|^{2}_{\infty}\leq dr\|\bm{L}_{1,j}-\bm{L}_{2,j}\|^{2}_{\mathrm{\infty}}$, and 3) $\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{F}}\leq\|\bm{f}_{1}-\bm{f}_{2}\|^{2}_{F}+2\|\bm{f}_{1}\|_{\infty}\|\bm{E}-\bm{E}_{0}\|_{F}^{2}\leq d\|\bm{f}_{1}-\bm{f}_{2}\|^{2}_{\infty}+2\|\bm{f}_{1}\|_{\infty}\|\bm{E}_{1}\|_{0}\|\bm{E}_{1}-\bm{E}_{2}\|_{\infty}^{2}$.
Using 1), 2) and 3) above, we have i) $\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\infty},\|\bm{L}_{1,j}-\bm{L}_{2,j}\|^{2}_{\infty}\leq\zeta_{T}^{2}/dr,$ ii) $\|\bm{f}_{1}-\bm{f}_{2}\|^{2}_{\infty}\leq\zeta_{T}^{2}/d,$ and iii) $\|\bm{E}_{1}-\bm{E}_{2}\|_{\infty}^{2}\leq\max\{({T}\bar{\epsilon}_{T}^{2})^{-2/\gamma}\log d,1/d^{2}\}\zeta_{T}^{2}$ implies $\{\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{F}},\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{F}},\|\bm{L}_{1,k}-\bm{L}_{2,k}\|^{2}_{\mathrm{F}}\}\lesssim\zeta_{T}^{2}$. Using the result from the above display, we let $\zeta_{T}^{2}$ as $\delta_{T}^{2}T^{-3}(T\bar{\epsilon}_{T}^{2})^{-2/\gamma}d^{-3-p}(T\bar{\epsilon}_{T}^{2})^{-(8+8p)/\gamma}$ Thus, at $\bm{\zeta}_{2}\in\mathcal{F}_{T}$ with $\|\bm{\Upsilon}_{T,2}-\bm{\Upsilon}_{T,1}\|_{\mathrm{F}}<\delta_{T}$, ${\mathbb{E}}_{\bm{\zeta}_{2}}(1-\phi_{T})\leq\exp[-c_{3}T\epsilon_{T}^{2}]$.
@@ A.3 Rest of the proof of Theorem 2
$\leq pC_{0}e^{-c_{0}C^{\gamma}T\bar{\epsilon}_{T}^{2}}+2pdrC_{0}e^{-c_{0}C^{\gamma}T\bar{\epsilon}_{T}^{2}}$
Since $\log d\ll T\bar{\epsilon}_{T}^{2}$, the above expression is bounded by $B_{0}e^{-b_{0}T\epsilon_{T}^{2}}$ for some $B_{0},b_{0}>0$ and $b_{0}$ can be chosen as large as we please by making $C$ in the definition of the sieve $\mathcal{F}_{T}$ large enough. Hence, by arguing as in the proof of Theorem 8.19 of Ghosal and van der Vaart [2017], it follows that the posterior contraction rate for $\bm{\Upsilon}_{T}$ at $\bm{\Upsilon}_{T,0}$ in terms of the Frobenius distance is $\sqrt{T}\sqrt{d^{2}(\log T)/T}=d\sqrt{(\log T)}$.
Let the collection of such $\bm{\zeta}_{2}$ with $\|\bm{\Upsilon}_{T,2}-\bm{\Upsilon}_{T,1}\|_{\mathrm{F}}<\delta_{T}$ be denoted by $\mathcal{B}_{T}=\{\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\infty},\|\bm{L}_{1,j}-\bm{L}_{2,j}\|^{2}_{\infty}\leq\zeta_{T}^{2}/dr,\|\bm{f}_{1}-\bm{f}_{2}\|^{2}_{\infty}\leq\zeta_{T}^{2}/d,\|\bm{E}_{1}-\bm{E}_{2}\|_{\infty}^{2}\leq\zeta_{T}^{2}/d^{2}\}$, where $\zeta_{T}^{2}=\delta_{T}^{2}T^{-3}(T\bar{\epsilon}_{T}^{2})^{-2/\gamma}d^{-3-p}(T\bar{\epsilon}_{T}^{2})^{-(8+8p)/\gamma}$. The number of such sets $\mathcal{B}_{T}$ needed to cover $\mathcal{F}_{T}$ is at most $N_{T}\leq T^{B}$ following the similar lines of arguments as in Lemma 8 of Roy et al. [2024] setting $\delta_{T}$ a negative polynomial in $T$. Therefore $\log N_{T}\lesssim T\epsilon_{T}^{2}$. For each $\mathcal{B}_{T}$ containing a point $\bm{\zeta}_{1}$ such that $\|\bm{\Upsilon}_{T,1}-\bm{\Upsilon}_{T,0}\|_{\mathrm{F}}>\epsilon_{T}$, we choose a representative and construct the likelihood ratio test $\phi_{T}$. The final test for testing the hypothesis $\bm{\zeta}=\bm{\zeta}_{0}$ against the alternative $\{\bm{\zeta}:\|\bm{\Upsilon}_{T,1}-\bm{\Upsilon}_{T,0}\|_{\mathrm{F}}>\sqrt{T}\epsilon_{T}\}$ is the maximum of these tests. Then, the resulting test has the probability of type II error bounded by $\exp[-c_{3}T\epsilon_{T}^{2}]$ while the probability of type I error is bounded by $\exp[\log N_{T}-c_{2}^{\prime}T\epsilon_{T}^{2}]\leq\exp[-c_{2}^{\prime}T\epsilon_{T}^{2}/2]$ if we choose $M>0$ sufficiently large.
@@ A.4 Proof of Theorem 3
When $\bm{\Omega}_{0}$ is sparse and the prior for $\bm{\Omega}$ imposes sparsity as described in Section 3.4, we show the improved rate $\sqrt{(d+s)(\log T)/T}$. To this end, we refine the estimate of the prior concentration in Subsection A.1. The arguments used for the test construction, bounding the prior probability of the complement of the sieve and for converting the rate result in terms of the Frobenius distance on $\bm{\Omega}$ and $\bm{R}_{1},\ldots,\bm{R}_{p}$ remain the same. Except to control the number of nonzero entries in $\bm{E}$, we add $\|\bm{E}\|_{0}\leq C({T}\bar{\epsilon}_{T}^{2})^{1/\gamma}/\log d$ in the sieve and let $\mathcal{B}_{T}=\{\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\infty},\|\bm{L}_{1,j}-\bm{L}_{2,j}\|^{2}_{\infty}\leq\zeta_{T}^{2}/dr,\|\bm{f}_{1}-\bm{f}_{2}\|^{2}_{\infty}\leq\zeta_{T}^{2}/d,\|\bm{E}_{1}-\bm{E}_{2}\|_{\infty}^{2}\leq({T}\bar{\epsilon}_{T}^{2})^{-2/\gamma}\log d\zeta_{T}^{2}\}$. Covering number calculation will remain identical to Lemma 8 of Roy et al. [2024].
Since $r$ is assumed to be a fixed constant, $-\log\Pi(\|\bm{L}_{j}-\bm{L}_{j,0}\|\leq\eta)\lesssim d\log(1/\eta)$, $j=1,\ldots,p$, and so is the corresponding estimate for $\bm{K}_{1},\ldots,\bm{K}_{p}$. For $\bm{f}$, the corresponding estimate $d\log(1/\eta)$ does not change from the previous scenario. However, $\bm{E}_{0}$ is now $s$-dimensional instead of $\binom{d}{2}$. The estimate for $-\log\Pi(\|\bm{E}-\bm{E}_{0}\|_{\infty}\leq\eta)$ under the shrinkage prior improves to a multiple of $(d+s)\log(1/\eta)$ by the arguments given in the proof of Theorem 4.2 of Shi et al. [2021]. Thus $-\log\Pi(\|\bm{\zeta}-\bm{\zeta}_{0}\|_{\infty}\leq\eta)\lesssim(d+s)\log(1/\eta)$, leading to the asserted pre-rate $\bar{\eta}=\sqrt{(d+s)(\log T)/T}$. We also note that the additional sparsity imposed on $\bm{L}_{1},\ldots,\bm{L}_{p}$ through the prior does not further improve the rate when $r$ is a fixed constant and it may be replaced by a fully non-singular prior.
@@
We have $\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{F}$ is small if $\frac{1}{T}\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{F}$ is small.
@@
If $\frac{1}{T}\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{F}$ is small, $\|\bm{A}_{1}-\bm{A}_{2}\|^{2}_{F}$ is small, where $\bm{A}$ is the autoregressive coefficient.
@@
where $M_{P,1}=[(p+1)+p^{3}\{M_{1,U}M_{1,V}M_{2,U}M_{2,V}+2C_{T}(M_{1,V}+M_{2,U})\}+2G_{T}\{p^{4}M_{1,U}M_{1,V}M_{1,C,U,V}+2p^{2}C_{T}M_{2,V}M_{1,V}M_{1,C,U,V}+2p^{2}C_{T}M_{2,U}M_{1,V}M_{1,C,U,V}\}]$ and $M_{P,2}=[2p^{2}(M_{1,U}M_{1,V}M_{1,C,U,V}+p)+2pC_{T}M_{2,V}M_{1,V}M_{1,C,U,V}+2pC_{T}M_{2,U}M_{1,V}M_{1,C,U,V}]$.
The proof is based on next 5 Lemmas and results.

### v2

@@ A.2 Test construction
Let $\rho_{1},\ldots,\rho_{p}$ stand for the eigenvalues of $\bm{\Upsilon}_{T,0}^{-1/2}(\bm{\Upsilon}_{T,1}-\bm{\Upsilon}_{T,0})\bm{\Upsilon}_{T,0}^{-1/2}$. By Lemma 4, the Reyni divergence is given by $\frac{1}{4}\sum_{j=1}^{T}[2\log(1+\rho_{j})-\log(1+\rho_{j})]$. By arguing as in the proof of Lemma 1 of Roy et al. (2024), each term inside the sum can be bounded below by $c_{2}\min(\rho_{j}^{2},c_{1})$ for some constants $c_{1},c_{2}>0$, where $c_{1}$ can be chosen as large as we like at the expense of making $c_{2}$ appropriately smaller. Under Condition (A4), using Relation (iii) of Lemma 5, we obtain that the $\rho_{j}$ are bounded by some constant not changing with $T$. Thus, with a sufficiently large $c_{1}$, the minimum operation in the estimate is redundant; that is, the Reyni divergence is bounded below by $\frac{1}{4}c_{2}\|\bm{\Upsilon}_{T,0}^{-1/2}(\bm{\Upsilon}_{T,1}-\bm{\Upsilon}_{T,0})\bm{\Upsilon}_{T,0}^{-1/2}\|_{\mathrm{F}}^{2}\geq c_{2}^{\prime}T\epsilon_{T}^{2}$ for some constant $c_{2}^{\prime}>0$. This follows since $\|\bm{\Upsilon}_{T,0}^{-1/2}(\bm{\Upsilon}_{T,1}-\bm{\Upsilon}_{T,0})\bm{\Upsilon}_{T,0}^{-1/2}\|_{\mathrm{F}}\geq\|\bm{\Upsilon}_{T,1}-\bm{\Upsilon}_{T,0}\|_{\mathrm{F}}/\|\bm{\Upsilon}_{T,0}\|^{2}_{\mathrm{op}}$ and by Relation (iii) of Lemma 5, $\|\bm{\Upsilon}_{T,0}\|^{2}_{\mathrm{op}}\leq\|\bm{\Omega}^{-1}_{0}\|^{2}_{\mathrm{op}}$, which is bounded by a constant by Condition (A5). Thus both error probabilities are bounded by $e^{-c_{2}^{\prime}T\epsilon_{T}^{2}}$ for some $c_{2}^{\prime}>0$.
@@
${\mathbb{E}}_{\bm{\zeta}_{2}}(1-\phi_{T})$ $\leq[{\mathbb{E}}_{\bm{\zeta}_{1}}(1-\phi_{T})]^{1/2}[{\mathbb{E}}_{\bm{\zeta}_{1}}(Q_{\bm{\zeta}_{2}}/Q_{\bm{\zeta}_{1}})^{2}]^{1/2}$
@@
which is bounded by
$\|\bm{f}_{1}\|_{\infty}+pdr\max_{j}(\|\bm{L}_{j,1}\|_{\infty}^{2})\leq C({T}\bar{\epsilon}_{T}^{2})^{1/\gamma}+pdr(T\bar{\epsilon}_{T}^{2})^{2/\gamma}\leq C^{\prime}dr(T\bar{\epsilon}_{T}^{2})^{2/\gamma}$
for some constant $C^{\prime}>0$ as $p$ is fixed and $T\bar{\epsilon}_{T}^{2}\to\infty$. Thus the expression in (A.2) is bounded by $\exp[-c_{2}T\epsilon_{T}^{2}+(C^{\prime}dr)^{2}(T\bar{\epsilon}_{T}^{2})^{4/\gamma}\delta_{T}^{2}]\leq\exp[-c_{3}T\epsilon_{T}^{2}]$ for some constant $c_{3}>0$ by choosing $\delta_{T}$ a sufficiently small constant multiple of $(C^{\prime}dr)^{-1}(T\bar{\epsilon}_{T}^{2})^{1/2-2/\gamma}$ and $M>0$ large enough.
@@
when $\max(\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{F}},\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{F}},\|\bm{L}_{1,k}-\bm{L}_{2,k}\|^{2}_{\mathrm{F}})\lesssim\zeta_{T}^{2}$
We have,
  • (1)
    $\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{F}}\leq dr\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\infty}$;
  • (2)
    $\|\bm{L}_{1,j}-\bm{L}_{2,j}\|^{2}_{\infty}\leq dr\|\bm{L}_{1,j}-\bm{L}_{2,j}\|^{2}_{\mathrm{\infty}}$;
  • (3)
    $\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{F}}$ is bounded by
    $\|\bm{f}_{1}-\bm{f}_{2}\|^{2}_{\mathrm{F}}+2\|\bm{f}_{1}\|_{\infty}\|\bm{E}-\bm{E}_{0}\|_{\mathrm{F}}^{2}\leq d\|\bm{f}_{1}-\bm{f}_{2}\|^{2}_{\infty}+2\|\bm{f}_{1}\|_{\infty}\|\bm{E}_{1}\|_{0}\|\bm{E}_{1}-\bm{E}_{2}\|_{\infty}^{2}.$
Using (1), (2), and (3) above, we have
  • (i)
    $\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\infty},\|\bm{L}_{1,j}-\bm{L}_{2,j}\|^{2}_{\infty}\leq\zeta_{T}^{2}/dr$;
  • (ii)
    $\|\bm{f}_{1}-\bm{f}_{2}\|^{2}_{\infty}\leq\zeta_{T}^{2}/d$;
  • (iii)
    $\|\bm{E}_{1}-\bm{E}_{2}\|_{\infty}^{2}\leq\max\{({T}\bar{\epsilon}_{T}^{2})^{-2/\gamma}\log d,1/d^{2}\}\zeta_{T}^{2}$ implies
    $\{\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{F}},\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{F}},\|\bm{L}_{1,k}-\bm{L}_{2,k}\|^{2}_{\mathrm{F}}\}\lesssim\zeta_{T}^{2}.$
Using the result from the above display, we let
$\zeta_{T}^{2}=\delta_{T}^{2}T^{-3}(T\bar{\epsilon}_{T}^{2})^{-2/\gamma}d^{-3-p}(T\bar{\epsilon}_{T}^{2})^{-(8+8p)/\gamma}.$
Thus, at $\bm{\zeta}_{2}\in\mathcal{F}_{T}$ with $\|\bm{\Upsilon}_{T,2}-\bm{\Upsilon}_{T,1}\|_{\mathrm{F}}<\delta_{T}$,
${\mathbb{E}}_{\bm{\zeta}_{2}}(1-\phi_{T})\leq\exp[-c_{3}T\epsilon_{T}^{2}].$
@@ A.3 Rest of the proof of Theorem 2
$\leq pC_{0}e^{-c_{0}C^{\gamma}T\bar{\epsilon}_{T}^{2}}+2pdrC_{0}e^{-c_{0}C^{\gamma}T\bar{\epsilon}_{T}^{2}}.$
Since $\log d\ll T\bar{\epsilon}_{T}^{2}$, the above expression is bounded by $B_{0}e^{-b_{0}T\epsilon_{T}^{2}}$ for some $B_{0},b_{0}>0$ and $b_{0}$ can be chosen as large as we please by making $C$ in the definition of the sieve $\mathcal{F}_{T}$ large enough. Hence, by arguing as in the proof of Theorem 8.19 of Ghosal and van der Vaart (2017), it follows that the posterior contraction rate for $\bm{\Upsilon}_{T}$ at $\bm{\Upsilon}_{T,0}$ in terms of the Frobenius distance is $\sqrt{T}\sqrt{d^{2}(\log T)/T}=d\sqrt{\log T}$.
Let the collection of such $\bm{\zeta}_{2}$ with $\|\bm{\Upsilon}_{T,2}-\bm{\Upsilon}_{T,1}\|_{\mathrm{F}}<\delta_{T}$ be denoted by
$\mathcal{B}_{T}$ =$=\{\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\infty},\|\bm{L}_{1,j}-\bm{L}_{2,j}\|^{2}_{\infty}\leq\zeta_{T}^{2}/dr,$
$\qquad\qquad\|\bm{f}_{1}-\bm{f}_{2}\|^{2}_{\infty}\leq\zeta_{T}^{2}/d,\|\bm{E}_{1}-\bm{E}_{2}\|_{\infty}^{2}\leq\zeta_{T}^{2}/d^{2}\},$
where $\zeta_{T}^{2}=\delta_{T}^{2}T^{-3}(T\bar{\epsilon}_{T}^{2})^{-2/\gamma}d^{-3-p}(T\bar{\epsilon}_{T}^{2})^{-(8+8p)/\gamma}$. The number of such sets $\mathcal{B}_{T}$ needed to cover $\mathcal{F}_{T}$ is at most $N_{T}\leq T^{B}$ following the similar lines of arguments as in Lemma 8 of Roy et al. (2024) setting $\delta_{T}$ a negative polynomial in $T$. Therefore $\log N_{T}\lesssim T\epsilon_{T}^{2}$. For each $\mathcal{B}_{T}$ containing a point $\bm{\zeta}_{1}$ such that $\|\bm{\Upsilon}_{T,1}-\bm{\Upsilon}_{T,0}\|_{\mathrm{F}}>\epsilon_{T}$, we choose a representative and construct the likelihood ratio test $\phi_{T}$. The final test for testing the hypothesis $\bm{\zeta}=\bm{\zeta}_{0}$ against the alternative $\{\bm{\zeta}:\|\bm{\Upsilon}_{T,1}-\bm{\Upsilon}_{T,0}\|_{\mathrm{F}}>\sqrt{T}\epsilon_{T}\}$ is the maximum of these tests. Then, the resulting test has the probability of type II error bounded by $\exp[-c_{3}T\epsilon_{T}^{2}]$ while the probability of type I error is bounded by $\exp[\log N_{T}-c_{2}^{\prime}T\epsilon_{T}^{2}]\leq\exp[-c_{2}^{\prime}T\epsilon_{T}^{2}/2]$ if we choose $M>0$ sufficiently large.
@@ A.4 Proof of Theorem 3
When $\bm{\Omega}_{0}$ is sparse and the prior for $\bm{\Omega}$ imposes sparsity as described in Section 3.4, we show the improved rate $\sqrt{(d+s)(\log T)/T}$. To this end, we refine the estimate of the prior concentration in Subsection A.1. The arguments used for the test construction, bounding the prior probability of the complement of the sieve and for converting the rate result in terms of the Frobenius distance on $\bm{\Omega}$ and $\bm{R}_{1},\ldots,\bm{R}_{p}$ remain the same. Except to control the number of nonzero entries in $\bm{E}$, we add $\|\bm{E}\|_{0}\leq C({T}\bar{\epsilon}_{T}^{2})^{1/\gamma}/\log d$ in the sieve and let
$\mathcal{B}_{T}$ =$=\{\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\infty}\leq\zeta_{T}^{2}/dr,\|\bm{L}_{1,j}-\bm{L}_{2,j}\|^{2}_{\infty}\leq\zeta_{T}^{2}/dr,$
$\qquad\|\bm{f}_{1}-\bm{f}_{2}\|^{2}_{\infty}\leq\zeta_{T}^{2}/d,\|\bm{E}_{1}-\bm{E}_{2}\|_{\infty}^{2}\leq({T}\bar{\epsilon}_{T}^{2})^{-2/\gamma}\log d\zeta_{T}^{2}\}.$
Covering number calculation will remain identical to Lemma 8 of Roy et al. (2024).
Since $r$ is assumed to be a fixed constant, $-\log\Pi(\|\bm{L}_{j}-\bm{L}_{j,0}\|\leq\eta)\lesssim d\log(1/\eta)$, $j=1,\ldots,p$, and so is the corresponding estimate for $\bm{K}_{1},\ldots,\bm{K}_{p}$. For $\bm{f}$, the corresponding estimate $d\log(1/\eta)$ does not change from the previous scenario. However, $\bm{E}_{0}$ is now $s$-dimensional instead of $\binom{d}{2}$. The estimate for $-\log\Pi(\|\bm{E}-\bm{E}_{0}\|_{\infty}\leq\eta)$ under the shrinkage prior improves to a multiple of $(d+s)\log(1/\eta)$ by the arguments given in the proof of Theorem 4.2 of Shi et al. (2021). Thus
$-\log\Pi(\|\bm{\zeta}-\bm{\zeta}_{0}\|_{\infty}\leq\eta)\lesssim(d+s)\log(1/\eta),$
leading to the asserted pre-rate $\bar{\eta}=\sqrt{(d+s)(\log T)/T}$. We also note that the additional sparsity imposed on $\bm{L}_{1},\ldots,\bm{L}_{p}$ through the prior does not further improve the rate when $r$ is a fixed constant and it may be replaced by a fully non-singular prior.
@@
We have $\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{F}}$ is small if $T^{-1}\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{\mathrm{F}}$ is small.
@@
If $T^{-1}\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{\mathrm{F}}$ is small, $\|\bm{A}_{1}-\bm{A}_{2}\|^{2}_{\mathrm{F}}$ is small, where $\bm{A}$ is the autoregressive coefficient.
@@
where $M_{P,1}$ is
$(p+1)+p^{3}\{M_{1,U}M_{1,V}M_{2,U}M_{2,V}+2C_{T}(M_{1,V}+M_{2,U})\}$
$+2G_{T}p^{2}\{p^{2}M_{1,U}M_{1,V}M_{1,C,U,V}+2C_{T}M_{2,V}M_{1,V}M_{1,C,U,V}+2C_{T}M_{2,U}M_{1,V}M_{1,C,U,V}\}$
and $M_{P,2}$ is
$2p^{2}(M_{1,U}M_{1,V}M_{1,C,U,V}+p)+2pC_{T}M_{2,V}M_{1,V}M_{1,C,U,V}+2pC_{T}M_{2,U}M_{1,V}M_{1,C,U,V}.$
The proof is based on the next five lemmas and results.